PURE ROTATIONAL SPECTROSCOPY OF ASYMMETRIC TOPS IN THE UNDERGRADUATE CLASS-ROOM OR LABORATORY

A. J. MINEI, Department of Chemistry and Biochemistry, Division of Natural Sciences, College of Mount Saint Vincent, 6301 Riverdale Avenue, Riverdale, New York, 10471; \underline{S. A. COOKE}, School of Natural and Social Sciences, Purchase College SUNY, 735 Anderson Hill Road, Purchase, NY 10577, USA.

Due to concerns of complexity, the asymmetric top, for which κ = (2B − A − C)/(A − C) ≠ ± 1, is feared, or at least avoided, by many instructors when explaining the rigid rotor. However, the spectral patterns formed by cold asymmetric rigid rotors in the centimeter-wave region of the electromagnetic spectrum can be easily identified. We will present some techniques for spectral analyses that we have successfully employed with undergraduate students who are either “prequantum mechanics” or are currently enrolled in a chemical quantum mechanics class. The activities are simple, requiring the students to first locate repeating patterns and then apply simple algebraic expressions in order to determine all three rotational constants. The method will be illustrated using the spectra of 2,2,3,3-tetrafluoropropyl trifluoroacetate (CF_{3}C(=O)OCH_{2}CF_{2}CHF_{2}), (E)-1,3,3,3-tetrafluoropropene (CF_{3}CH=CHF), 1H,1H,2H-perfluorocyclobutane (CF_{2}CF_{2}CHFCH_{2}), and 2H-nonafluorobutane (CF_{3}CHFCF_{2}CF_{3}). The first two of these species have predominantly a-type spectra, the third has a predominantly b-type spectrum, the fourth has a predominantly c-type spectrum.